OpenStudy (anonymous):

university 1st year math question.

OpenStudy (welshfella):

i looked up torricelli's law and only found a formula for a cylindrical tank..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i am finding it very difficult

ganeshie8 (ganeshie8):

@ParthKohli try this

Parth (parthkohli):

are you mathmath333?

OpenStudy (anonymous):

what ?

ganeshie8 (ganeshie8):

No, he is not

Parth (parthkohli):

I haven't studied fluids yet, but I looked at the Wikipedia page and it sorta makes sense.

Parth (parthkohli):

This is what we're really looking at though. https://en.wikipedia.org/wiki/Torricelli%27s_law#Application_for_time_to_empty_the_container

ganeshie8 (ganeshie8):

same here, it seems torcelli's law gives the speed of fluid of height \(h\), leaving tank from bottom : \[v = \sqrt{2gh}\]

OpenStudy (anonymous):

noo \[v=\sqrt{2gh}\]

OpenStudy (ytrewqmiswi):

What does that k represents?

OpenStudy (anonymous):

i was thinking of an approach where i find A(h) and use integration

Parth (parthkohli):

Is that the area of cross-section as a function of height? And are you sure this is the equation?\[A(h) \frac{dh}{dt} = -k \sqrt{h}\]

OpenStudy (anonymous):

yes 100% because we havent used other v=sqrt2gh yet and my teacher wouldnt expect us to use that

OpenStudy (anonymous):

the only problem i am having is how to find A(h)

OpenStudy (ytrewqmiswi):

That wuld mean that the area of cross section varies with height will that make any sense?

Parth (parthkohli):

But in your equation, \(\sqrt{2g}\) or \(g\) won't appear. We can find the area of the cross-section as a function of height.

Parth (parthkohli):

If my eyes work, we're looking at a rectangular cross-section, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and yes i wan to find the area as a function of height plz

Parth (parthkohli):

Width increases in a linear manner from 4 to 5. At h = 0, it is 4 and at h = 4, it is 5. Thus in general it should be \(w(h) = h/4 + 4\). Meanwhile the length remains 10 so \(l(h) = 10\). Multiply...

OpenStudy (anonymous):

\[\frac{ 10h }{ 8 }\]

Parth (parthkohli):

noooo I didn't write h/8 h/4+4 means (1/4)*h + 4

OpenStudy (anonymous):

\[\frac{ 10(h+16) }{ 4 }\]

ganeshie8 (ganeshie8):

40----> 50 A(y) = 40 + (y/4)*10 does that work ?

Parth (parthkohli):

looks good

Parth (parthkohli):

it only works because one of them remains constant can't do the same thing when both width and length increase which is why it's a good practice to do them separately

ganeshie8 (ganeshie8):

yeah 10cm is fixed, only the other dimension is changing, so it is linear

OpenStudy (anonymous):

just a question srry interpreting ur thought process how did u get width as (1/4) times h +4

ganeshie8 (ganeshie8):

\[A(h)\frac{ dh }{ dt }=-k \sqrt{h}\] plugin \(A(h)\), the de becomes \[(40+(\frac{h}{4})*10)\frac{ dh }{ dt }=-k \sqrt{h}\]

OpenStudy (ytrewqmiswi):

the liquid escapes with a velocity =sqrt(2gh) which varies with the level of liquid. Hence, we have to use integral. Let y be the height of the liquid at an instant. This height changes by dy in time dt. Volume of water leaving out per secound=-Ady.dt at the hole volume escaping per second is av=a[sqrt(2gy)] a[sqrt(2gy)] =-Ady/dt ---->\[\int\limits_{H}^{0}\frac{ -dy }{ \sqrt{y} } =\frac{ a \sqrt{2g} }{ A }\int\limits_{t}^{0}dt\] \[t=\frac{ A }{ a \sqrt{2g}}2\sqrt{H}\] a<<A? what did i do o.o

ganeshie8 (ganeshie8):

maybe think of it in reverse : A(h) = 40 + (h/4)*10 plugin h=0, what do you get ? plugin h=4, what do you get ?

Parth (parthkohli):